If I rent a camera for $128 for 4 days, how much would it coust to rent for 6 months?
Answer by srinivas.g(540)

rent for 4 days =$128
rent for 1 day = 128/4
= $32
6 months =6*30
= 180 days
rent for 180 days = 180*32
= $5760
Rent = $5760